# Dévissage for generation in derived categories

### v4

@@ Abstract.
This note is concerned with generation in the derived category of bounded complexes with coherent cohomology over a Noetherian scheme. We demonstrate a flavor of ‘dévissage’ by identifying two explicit collections of structure sheaves for closed subschemes that classically generate the bounded derived category. Amongst the two, one consists of those supported on the singular locus of the scheme. Moreover, building from the work of Aoki, we show the essential image of the derived pushforward along a proper surjective morphism of Noetherian schemes strongly generates the targets bounded derived category.
@@ 1. Introduction
Recall a notion of generation for a triangulated category $\mathcal{T}$, which was introduced in [BvdB03]. A subcategory $\mathcal{S}$ in $\mathcal{T}$ is said to classically generate if the smallest thick subcategory containing $\mathcal{S}$, denoted $\langle\mathcal{S}\rangle$, coincides with $\mathcal{T}$. This means every object in $\mathcal{T}$ can be obtained from objects in $\mathcal{S}$ using cones, shifts, finite coproducts, and retracts. Denote by $\langle\mathcal{S}\rangle_{n}$ for the smallest full subcategory of $\mathcal{T}$ generated from $\mathcal{S}$ using finite coproducts, shifts, retracts, and at most $n$ cones. We will say $\mathcal{S}$ strongly generates if $\langle\mathcal{S}\rangle_{n}=\mathcal{T}$ for some integer $n\geq 0$. In the special case $\mathcal{S}$ consists of a single object $G$, then we write $\langle\mathcal{S}\rangle$ as $\langle G\rangle$ and $\langle\mathcal{S}\rangle_{n}$ as $\langle G\rangle_{n}$.
The concept of generators has led to interesting invariants, as introduced in [ABIM10, Rou08]. There are important benefits a triangulated category enjoys in the presence of a single object which strongly generates, including representability results on cohomological functors [Rou08]. Those that are geometrically inclined might appreciate a connection between generation and openness of the regular locus for a Noetherian scheme [IT19, DL24].
We will focus on the derived category of bounded complexes with coherent cohomology over a Noetherian scheme $X$, which is denoted by $D^{b}_{\operatorname{coh}}(X)$. There has been considerable progress made in understanding instances where $D^{b}_{\operatorname{coh}}(X)$ can be generated by a single object. These cases have been studied for quasi-excellent separated schemes of finite Krull dimension [Aok21, Nee21], Noetherian $J\textrm{-}2$ schemes [ELS20], Noetherian schemes admitting a separator [Jat21], and various classes of affine Noetherian schemes [IT16, IT19, DLT23]. There have been efforts to explicitly identify such objects [Rou08, BIL+23], including noncommutative techniques [BDL23, BDG17, DLR24].
Our motivation is classical at heart. An intermediate step in arguments for generic freeness à la dévissage asserts that for any finitely generated module $M$ over a commutative Noetherian ring $R$ there exists a finite descending chain of submodules $\cdots\subseteq M_{1}\subseteq M_{0}=:M$ whose corresponding quotients $M_{i}/M_{i+1}$ are isomorphic to objects of the form $R/\mathfrak{p}$ for some $\mathfrak{p}$ is in $\operatorname{Spec}(R)$. This is observed in the proof of [Gro65, Lemme 6.9.2]. If interpreted in our context, this tells us that any finitely generated $R$-module can be obtained from the collection of modules of the form $R/\mathfrak{p}$ where $\mathfrak{p}$ is in $\operatorname{Spec}(R)$ using cones, shifts, and direct sums in $D^{b}_{\operatorname{coh}}(\operatorname{Spec}(R))$.
This brings attention to our first result where we demonstrate a flavor of ‘dévissage’ for derived categories associated to schemes.
Theorem A.
(see Theorem 3.3) Let $X$ be a Noetherian scheme. If $P$ is a perfect complex on $X$ with full support, then $D_{\operatorname{coh}}^{b}(X)$ is classically generated by $P$ and objects of the form $i_{\ast}\mathbb{L}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme.
@@
There is an interesting application of Theorem A when applied to the special case of $P=\mathcal{O}_{X}$. This tells us that any object in $D^{b}_{\operatorname{coh}}(X)$ can be obtained, using finitely many cones, from $\mathcal{O}_{X}$ and those of the form $i_{\ast}\mathcal{O}_{Z}$ for $i\colon Z\to X$ a closed immersion from an integral scheme. However, there is a refinement for Theorem A in the affine setting. Namely, [Tak14] tells us that $D^{b}({\operatorname{mod}}R)$ is classically generated by $R$ and $R/\mathfrak{p}$ for $\mathfrak{p}$ in the singular locus of $R$. Recall the singular locus of a Noetherian scheme $X$, denoted $\operatorname{sing}(X)$, is the collection of points $p$ in $X$ such that $\mathcal{O}_{X,p}$ is not regular.
Our next result tells us how this can globalized, where we leverage what is known affine locally.
Theorem B.
@@
(see Theorem 3.5) Let $X$ be a Noetherian scheme. If $P$ is a classical generator for $\operatorname{perf}(X)$, then $D^{b}_{\operatorname{coh}}(X)$ is classically generated by $P$ and the objects of the form $i_{\ast}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
The utility for Theorem B is demonstrated when the singular locus is finite as it allows for one to explicitly identify generators, see Example 3.7 for case of normal surfaces. We would like to note that similar statements have been observed for special cases where one has even further control using codimension of the closed subschemes, cf. [MS23, Lemma 2.9] and [Lan24, Proposition 4.7]. Theorem B has been under various guises in the literature for special settings, see Remark 3.6 for details.
A common method to detecting whether or not $D^{b}_{\operatorname{coh}}(X)$ can be generated by a single object is typically done by studying the behavior of the derived pushforward. These strategies are connected to Theorem 3.3 in an interesting way: It is possible to show that $\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle=D^{b}_{\operatorname{coh}}(X)$ for $\pi\colon Y\to X$ a proper surjective morphism. By a different vein, [Aok21] demonstrated a general case where $D^{b}_{\operatorname{coh}}(X)$ can be strongly generated by a single object. This was accomplished using a notion of descendability for algebras in tensor triangulated categories [Mat16, BS17].
The following builds a bit further on [Aok21] to tell us how to strengthen the connection of Theorem 3.3 and studying generation through the derived pushforward.
Theorem C.
(see Theorem 3.11) If $\pi\colon Y\to X$ is a proper surjective morphism of Noetherian schemes, then there is an integer $n$ such that $D^{b}_{\operatorname{coh}}(X)=\langle\mathbb{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle_{n}$.
Proposition C shows that generators in the bounded derived category remain preserved under derived pushforward along proper surjective morphisms of Noetherian schemes, see Corollary 3.12. This aids in explicitly identifying strong generators in the bounded derived category of coherent sheaves for varieties over fields through alterations, see Example 3.13. However, the consideration of generation by a derived pushforward of a morphism of schemes is very natural, and has been studied in the past.
An instance this phenomenon occurs is for a variety $X$ with rational singularities over a field of characteristic zero. If $\pi\colon\widetilde{X}\to X$ is a resolution of singularities, then $\mathbb{R}\pi_{\ast}\colon D^{b}_{\operatorname{coh}}(\widetilde{X})\to D^{b}_{\operatorname{coh}}(X)$ is essentially dense, cf. [Har17, Theorem 1.1], [Fav09, Corollary 3.3.8], or [Kaw06, Lemma 7.4]. That is, any object $E$ in $D^{b}_{\operatorname{coh}}(X)$ is a retract of $\mathbb{R}\pi_{\ast}E^{\prime}$ for some $E^{\prime}$ in $D^{b}_{\operatorname{coh}}(\widetilde{X})$. Moreover, in the case of rational singularities, [MS23] observed that the essential image of $\mathbb{R}\pi_{\ast}$ generates $D^{b}_{\operatorname{coh}}(X)$ solely through iterated cones and shifts, without retracts, which holds important $K$-theoretic implications. However, the significance of Proposition C lies in its inclusion of retracts, which allows for a broader study of the bounded derived category for varieties beyond those with rational singularities.
On the other hand, if one is working in characteristic zero, then derived schemes are intimately related to what are called differential graded schemes, and so many our of results have an interesting interpretation in such context, cf. [Gai13, GR17a, GR17b] for details. In particular, Theorem 3.11 provides an alternative proof to related result in the literature in regards to proper descent for derived algebraic geometry over a field of characteristic zero, cf. [Gai13, Corollary 8.1.3].
Notation 1.1.
@@
    $\operatorname{perf}X$ is the strictly full subcategory of $D(X)$ consisting of perfect complexes
@@ 2. Generation
First, we will briefly cover basic notions on generation in triangulated categories. For a further treatment on necessary background, the reader is referred to [BvdB03, Rou08]. Let $\mathcal{T}$ be a triangulated category with shift functor $[1]\colon\mathcal{T}\to\mathcal{T}$.
@@
If $\mathcal{S}$ consists of a single object $G$, then we write $\langle G\rangle$ for $\langle\mathcal{S}\rangle$.
@@
Remark 2.4.
Let $\mathcal{S}$ be a subcategory of $\mathcal{T}$. It is easy to verify for any object in $\langle\mathcal{S}\rangle$ there exists an object $S$ in $\langle\mathcal{S}\rangle_{1}$ such that $S$ finitely builds $E$. We will freely utilize this fact in Section 3.
@@
    $\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i}$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ if $\mathcal{L}$ is an ample line bundle on a smooth quasi-projective variety $X$ over a field, see [Rou08, Proposition 7.9]
@@
    $F_{\ast}^{e}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ if $\mathcal{L}$ is an ample line bundle on a singular quasi-projective variety $X$ over a perfect field of positive characteristic where $F^{e}\colon X\to X$ is the $e$-th iterate of the Frobenius morphism on $X$, see [BIL+23, Corollary 3.9].

### v5

@@ Abstract.
We study a form of dévissage for generation in derived categories of Noetherian schemes. First, we extend a result of Takahashi from the affine context to the global setting, showing that the bounded derived category is classically generated by a perfect complex together with structure sheaves of closed subschemes supported on the singular locus. Second, we make an observation for how generation behaves under the derived pushforward of a proper surjective morphism between Noetherian schemes. These results enable us to explicitly identify strong generators for projective schemes with isolated singularities and for singular varieties over a perfect field.
@@ 1. Introduction
Our work focuses on a form of dévissage in the context of generation in derived categories of a Noetherian scheme. This is a process that allows for a better understanding of a collection of objects in such categories by studying a smaller subcollection. Firstly, we identify a subcategory that is constructed from a single perfect complex and the singular locus, which classically generates the bounded derived category of coherent sheaves. Secondly, we make a useful observation regarding the behavior of strong generation under the derived pushforward associated with a proper surjective morphism between Noetherian schemes.
@@
We remind ourselves on a notion of generation for a triangulated category $\mathcal{T}$ that was introduced in [BvdB03]. A subcategory $\mathcal{S}$ classically generates $\mathcal{T}$ if the smallest thick subcategory containing $\mathcal{S}$, denoted $\langle\mathcal{S}\rangle$, equals $\mathcal{T}$. That is, every object in $\mathcal{T}$ can be built from $\mathcal{S}$ using cones, shifts, finite coproducts, and retracts. For $n\geq 0$, $\langle\mathcal{S}\rangle_{n}$ denotes the smallest full subcategory generated from $\mathcal{S}$ using these operations with at most n1n-1 cones (see Definition 2.3 for details). The subcategory $\mathcal{S}$ strongly generates $\mathcal{T}$ if $\langle\mathcal{S}\rangle_{n}=\mathcal{T}$ for some $n$. When $\mathcal{S}$ consists of a single object $G$, we write $\langle G\rangle$ and $\langle G\rangle_{n}$. The Rouquier dimension of $\mathcal{T}$ is the smallest $n\geq 0$ such that $\langle G\rangle_{n+1}=\mathcal{T}$ for some $G\in\mathcal{T}$; it is set to $\infty$ otherwise.
The focus will be on the bounded derived category of coherent sheaves over a Noetherian scheme $X$, denoted $D^{b}_{\operatorname{coh}}(X)$. There is a deep connection between classical generation by a single object and the openness of the regular locus in Noetherian schemes [IT19, DL25]. This tells us the geometry can influence the presence of being generated by a single object. Recent work has developed our understanding of when $D^{b}_{\operatorname{coh}}(X)$ can be generated by a single object, with efforts using commutative and noncommutative techniques [^1] A far from comprehensive list includes: [Aok21, Nee21, ELS20, Jat21, IT16, IT19, DLT25, Rou08, BIL+23, BLD25, BDG17, DLR24]..
Despite these efforts, a comprehensive picture of generation in these derived categories remains a mystery. To address this, we shift gears by looking at subcategories of $D^{b}_{\operatorname{coh}}(X)$ which classically generate. The first candidate we had in mind comes from structure sheaves of closed subschemes.
From [Sta23, Tag 01YF], given a Noetherian scheme $X$, we can infer that every object $\operatorname{coh}(X)$ belongs to the thick subcategory $\mathcal{T}$ generated by $\mathcal{O}_{X}$ and $i_{\ast}\mathcal{O}_{Z}$ where $i\colon Z\to X$ is a closed immersion from an integral scheme [^2] See proof of [Gro65, Lemme 6.9.2] for variation in the affine case. There is the distinguished triangle for any $E\in D^{b}_{\operatorname{coh}}(X)$:
$Z(E)\to E\to B(E)[1]\to Z(E)[1]$
where $Z(E)$ and $B(E)$ denote the cycle and boundary complexes respectively. These complexes decompose into finite direct sums of shifts of coherent sheaves in $\operatorname{coh}(X)$. Then it follows that $\mathcal{T}=D^{b}_{\operatorname{coh}}(X)$. Moreover, one could tensor $\mathcal{T}$ with any perfect complex $P$ with full support and use projection formula to achieve the same outcome for classical generation. See Proposition 3.2 for an alternative argument.
@@
There is an interesting refinement for the observation above in the affine setting. Namely, [Tak14] tells us for $X$ an affine Noetherian scheme that $D^{b}_{\operatorname{coh}}(X)$ can be classically generated by $\mathcal{O}_{X}$ and structure sheaves of closed subscheme support in the singular locus of $X$. Recall the singular locus of $X$ is the collection of $p\in X$, denoted by $\operatorname{sing}(X)$, such that $\mathcal{O}_{X,p}$ is not a regular local ring.
The following result tells us how this can globalized.
Theorem A (see Theorem 3.3).
@@
Let $X$ be a Noetherian scheme. Suppose $P$ is a classical generator for $\operatorname{Perf}(X)$. Then $D^{b}_{\operatorname{coh}}(X)$ is classically generated by $P$ and the objects of the form $\mathbf{R}i_{\ast}\mathbf{L}i^{\ast}P$ where $i\colon Z\to X$ is a closed immersion from an integral scheme such that $i(Z)$ is contained in $\operatorname{sing}(X)$.
The utility for A is demonstrated when the singular locus is finite as it allows for one to explicitly identify generators. An exciting case is that for Noetherian scheme with isolated singularities (e.g. normal quasi-projective surfaces over a field) we have an explicit object that classically generates:
Corollary B.
Let $X$ be a scheme with isolated singularities $p_{1},\ldots,p_{n}$ that is projective over a Noetherian local ring $R$. Denote the closed immersion by $t\colon X\to\mathbb{P}^{N}_{R}$. Set $\mathcal{L}$ to be the line bundle $t^{\ast}\mathcal{O}_{\mathbb{P}^{n}_{R}}(-1)$. Then $D^{b}_{\operatorname{coh}}(X)$ is classically generated by the following object:
$(\bigoplus^{N}_{l=0}\mathcal{L}^{\otimes l})\oplus(\bigoplus^{n}_{j=1}i_{j,\ast}\mathcal{O}_{\operatorname{Spec}(\kappa(p_{j}))})$
where $i_{j}\colon\operatorname{Spec}(\kappa(p_{j}))\to X$ is the associated closed immersion for each $j$.
This is a special case of a more general setting for Noetherian schemes with isolated singularities (see Example 3.5). Observe its utility for projective varieties over a field as it identifies an explicit strong generator. To the best of our knowledge, the explicitness of such an object is very new even for mixed or positive characteristic settings. It becomes an interesting problem for future work to now leverage this concrete object in studying the Rouquier dimension of $D^{b}_{\operatorname{coh}}(X)$ for varieties $X$ over a field with isolated singularities.
A common method to detecting whether or not $D^{b}_{\operatorname{coh}}(X)$ can be generated by a single object is typically done by studying the behavior of the derived pushforward. Such strategies has appeared in the setting of varieties with rational singularities [Har17, Theorem 1.1], [Fav09, Corollary 3.3.8], or [Kaw06, Lemma 7.4].
Our next goal was to establish analogous results in a much broader context. Initially, we observed that it is possible to prove that $\langle\mathbf{R}\pi_{\ast}D^{b}_{\operatorname{coh}}(Y)\rangle=D^{b}_{\operatorname{coh}}(X)$ whenever $\pi\colon Y\to X$ is a proper surjective morphism. With the help of new key lemmas, we were then able to adapt an argument of Aoki [Aok21] to strengthen this result, obtaining a statement about strong generation in the derived category. Details can be found in Lemmas 3.6, 3.7, 3.8 and 3.9.
We push the upgrade further to say something about generation by a single object.
Proposition C (see Proposition 3.10).
Let $\pi\colon Y\to X$ be a proper surjective morphism of Noetherian schemes. If $G$ is a classical (resp. strong) generator for $D^{b}_{\operatorname{coh}}(Y)$, then $\mathbf{R}\pi_{\ast}G$ is a classical (resp. strong) generator for $D^{b}_{\operatorname{coh}}(X)$.
C gives generalizations of [Lan24, Theorem E], [LO25, Lemma 3.15], and [Gai13, Corollary 8.1.3]. A convenient consequence to C is that one can explicitly write down a strong generator for varieties over a perfect field.
Corollary D (see Example 3.11).
Let $X$ be a variety over a perfect field. By [dJ96, Theorem 4.1], there exists a proper surjective morphism $\pi\colon\widetilde{X}\to X$ with $\widetilde{X}$ is a regular quasi-projective variety. Let $\mathcal{L}$ be a very ample line bundle on $\widetilde{X}$. Then $D^{b}_{\operatorname{coh}}(X)$ is strongly generated by the object $\mathbf{R}\pi_{\ast}(\oplus^{\dim X}_{l=-\dim X}\mathcal{L}^{\otimes l})$.
Similar to B, we construct an explicit object that strongly generates the bounded derived category of coherent sheaves, $D^{b}_{\operatorname{coh}}(X)$, for any singular variety $X$ defined over a perfect field. We anticipate that this object will prove to be very useful in future work. Particularly, in the study of the Rouquier dimension for varieties with singularities that are not isolated (i.e. where B is not applicable).
Notation.
@@
    $\operatorname{Perf}(X)$ is the strictly full subcategory of $D(X)$ consisting of perfect complexes
@@ 2. Generation
We briefly cover generation in triangulated categories. For a further treatment on necessary background, the reader is referred to [BvdB03, Rou08]. Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$.
@@
If $\mathcal{S}$ consists of a single object $G$, then we write $\langle G\rangle$ for $\langle\mathcal{S}\rangle$. Additionally, we write $\mathcal{S}\star\mathcal{S}$ for the collection of $E\in\mathcal{T}$ which fit into a distinguished triangle
$S_{1}\to E\to S_{2}\to S_{1}[1]$
where $S_{1},S_{2}\in\mathcal{S}$. For any $n\geq 1$, we set
$\mathcal{S}^{\star n}:=\underbrace{\mathcal{S}\star\cdots\star\mathcal{S}}_{n\textrm{ times}}.$
@@
    $\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i}$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ if $\mathcal{L}$ is a very ample line bundle on a smooth quasi-projective variety $X$ over a field and $N\gg 0$, see [Orl09, Theorem 4].
@@
    $F_{\ast}^{e}(\bigoplus^{\dim X}_{i=0}\mathcal{L}^{\otimes i})$ is a strong generator for $D^{b}_{\operatorname{coh}}(X)$ if $\mathcal{L}$ is a very ample line bundle on a singular quasi-projective variety $X$ over a perfect field of positive characteristic where $F^{e}\colon X\to X$ is the $e$-th iterate of the Frobenius morphism on $X$, see [BIL+23, Corollary 3.9].